# Generalize across Homophily and Heterophily: Hybrid Spectral Graph Pre-Training and Prompt Tuning

### v2

@@ 3.2 Theoretical Analysis
Spectral Specificity in Pre-trained GNNs.
As the spectral distribution varies across graphs and feature dimensions, the optimal spectral characteristics of graph filters would also vary across graphs with different spectral distributions. Pretraining a fixed graph filter to adapt to different graphs would lead to poor generalization results.
To prove this, we focus on binary node classification since the multi-class classification can be simplified as multiple binary classifications (Chen, Ma, and Wang 2022). Let $\mathbf{Y}\in\mathbb{R}^{N\times 2}$ be the label matrix, with columns $\mathbf{y}_{0},\mathbf{y}_{1}$ as class indicators. Define label difference $\Delta\mathbf{y}=\mathbf{y}_{0}-\mathbf{y}_{1}$. Let $\hat{\mathbf{y}}=(\hat{y}_{1},\cdots,\hat{y}_{N})=\mathbf{U}^{T}\Delta\mathbf{y}$ and $\hat{\mathbf{x}}=(\hat{x}_{1},\cdots,\hat{x}_{N})$. Using Spectral Regression Loss (SRL) (Lei et al. 2022) for graph filter evaluation: $\mathcal{L}(\mathcal{G})=\sum_{i=1}^{N}\Big{(}\frac{\hat{y}_{i}}{\sqrt{N}}-\frac{g(\lambda_{i})\hat{x}_{i}}{\sqrt{\sum_{j=1}^{N}\big{(}{g(\lambda_{j})}^{2}\hat{x}_{j}^{2}\big{)}}}\Big{)}^{2}$. A smaller SRL indicates better downstream performance. With this, we propose Theorem 3.3.
@@
Given a pre-trained graph filter $g$, for an input graph $\mathcal{G}$ along with its spectral graph signal $\mathbf{\hat{x}}=(\hat{x}_{1},\hat{x}_{2},\ldots,\hat{x}_{N})^{T}$, in order to ensure a lower upper bound of the SRL, the spectral filter $g(\lambda_{i})$ should exhibit a positive correlation with $\hat{x}_{i}^{2}$. That is, a larger value of $g(\lambda_{i})$ ought to be assigned to a larger $\hat{x}_{i}^{2}$.
The complete proof is in Appendix A.2. Theorem 3.3 shows pre-trained graph filter specificity: a graph’s optimal filter should align spectrally with the graph itself, as optimal transfer requires alignment between pre-trained spectral filters and the downstream graph’s intrinsic spectrum. Thus, a single filter cannot generalize across graphs with varying spectral distributions.
While prior work addresses high-frequency importance for heterophilic graphs, we extend this insight to a more general conclusion: for graphs with varying homophily, abundant spectral knowledge (including not only low- and high-frequency but also band-pass mid-frequency information) is crucial. We verify the claim by testing spectral graph filters (i.e., low/mid/high-pass) on CSBM datasets from Figure 1. Results (details in Appendix B.5) show distinct graph filters excel on different graph types, aligning with Theorem 3.3.
@@ 4 Methodology
In Section 3.2, we empirically and theoretically verify that a single graph filter fails to generalize across graphs with varying spectral distributions. To address this issue, we first adopt hybrid spectral graph filters as our backbone (Section 4.1). During the pre-training stage, the model not only acquires knowledge for each graph filter, but also learns how to integrate representations from different graph filters (Section 4.2). In the prompt tuning stage, we propose prompt graphs to align the spectral distribution of downstream graphs with pretexts, facilitating the knowledge transfer (Section 4.3). The overall framework is in Figure 2, Algorithm 1, and Algorithm 2 in the appendix. Next, we give details.
@@ 4.1 Hybrid Spectral Filter Backbone
Theorem 3.3 shows that the filter performs well on graphs with similar spectral characteristics. Thus, to generalize across graphs with varying spectral distributions, we need to capture information from a wide range of frequencies. Recently, several works (Tang et al. 2022; Xu et al. 2024) attempt to utilize the Beta distribution as the basis for the graph filter, thereby obtaining the Beta wavelet GNN (BWGNN). This approach is favored for its excellent spectral locality, which allows BWGNN to consist of a set of filters, each focusing on different frequency components. Thus, we adopt BWGNN as our backbone model, and we also evaluate the performance of other hybrid GNN backbones in Section 5.4. The graph filters can be formulated as:
$g_{k,r}(\mathbf{L})=\frac{{\frac{\mathbf{L}}{2}}^{k}(\mathbf{I}-\frac{\mathbf{L}}{2})^{r}}{2B(k+1,r+1)},$ (6)
where $B(k+1,r+1)=\frac{k!r!}{(k+r+1)!}$ is a constant. With different constant values of $k$ and $r$, $g_{k,r}$ possesses different spectral characteristics. Setting $k+r=C$ as a constant, we obtain a group of $C+1$ graph filters $g_{\beta}=(g_{0,C},g_{1,C-1},\cdots,g_{C,0})$. Here, $g_{k,C-k}$ is a $C$ power polynomial, indicating $C$-hop neighbors are considered. Specifically, $g_{0,C}$ is a low-pass filter, $g_{C,0}$ is a high-pass filter, and others are band-pass filters of different frequencies. More details and visualization of BWGNN are provided in Appendix C.1 and Figure 6.
With the above hybrid graph filters, we can obtain a group of representations from different filters:
$\mathbf{Z}^{k}=MLP\Big{(}g_{k,C-k}(\mathbf{L})\mathbf{X}\Big{)}.$ (7)
@@ 4.2 Hybrid Spectral Pre-training
In Equation 7, graph filters generate representations seen as different spectral views. In pre-training, we train these filters using self-supervised tasks. Building on (Chen, Lei, and Wei 2024), which uses local-global contrastive learning (Velickovic et al. 2019) to align low- and high-pass views, we extend this to multiple spectral views. We perform contrastive learning between node embeddings in different views (as local patches) and integrated graph embeddings (as global summaries), as shown in the upper part of Figure 2.
@@
To obtain the global summaries, we first integrate node representations from different spectral views. Given the spectral distribution diversity in feature dimensions, we generate the integrated embedding for each node using an element-wise weighted sum:
$\mathbf{z}_{i}=\sum_{k=0}^{C}\frac{\exp(\mathbf{w}^{k})}{\sum_{j=0}^{C}\exp(\mathbf{w}^{j})}\odot\mathbf{z}_{i}^{k},$ (8)
@@
where $\odot$ denotes the element-wise product operation and $\mathbf{w}^{k}$ denotes the integration weight vector for $\mathbf{z}_{i}^{k}$. Then we perform mean pooling on the graph to get the global summary $\mathbf{z_{g}}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{z}_{i}$.
@@
To generate negative samples, we corrupt the graph structure by shuffling the nodes in the graph while keeping the edge index unchanged. Then we feed the corrupted graph into hybrid graph filers in Equation 7 to obtain negative samples $\mathbf{z}_{i}^{k-}$. We use a discriminator $\mathcal{D}$ to measure the agreement between node and graph embeddings $\mathcal{D}(\mathbf{z}_{i}^{k},\mathbf{z}_{g})=\sigma(\mathbf{z}_{i}^{k}\mathbf{W}\mathbf{z}_{g}^{T})\in[0,1]$, where $\mathbf{W}\in\mathbb{R}^{d\times d}$ denotes the weight matrix and $\sigma$ denotes the sigmoid activation function. With the above notations, the local-global agreement training objective can be formulated as:
$\mathcal{L}_{pre}(\mathbf{\theta})=-\frac{1}{CN}\sum_{k=0}^{C}\sum_{i=1}^{N}\Big{(}\log\mathcal{D}(\mathbf{z}_{i}^{k},\mathbf{z}_{g})+\log(1-\mathcal{D}(\mathbf{z}_{i}^{k-},\mathbf{z}_{g}))\Big{)}.$ (9)
The learnable parameters $\mathbf{\theta}$ include MLP parameters in each graph filter, integration weight vectors $\mathbf{w}^{k}$ for k=(0,1,,C)k=(0,1,...,C), and the weight matrix of discriminator $\mathcal{D}$. In this way, we can generate node representations with global information and maximize the mutual information between different spectral views (Velickovic et al. 2019; Chen, Lei, and Wei 2024). During pre-training, the model not only gains knowledge for each graph filter but also learns to integrate representations from different graph filters.
@@
Table 1: Performance comparison on 5-Shot node classification under the transductive setting. Here bold signifies the best result across all methods, while underline highlights the best baseline result.
@@
Table 2: Performance comparison on 5-Shot node classification under the inductive setting. Here ‘Source’ denotes the pre-training datasets and ‘Target’ denotes the downstream dataset.
4.3 Spectral-Aligned Prompt Tuning
Since graph filter excels in graphs with similar spectral distribution, in the prompt tuning stage, we aim to align the spectral distribution of downstream graphs with pretexts, thereby facilitating knowledge transfer. The process is shown in the lower part of Figure 2.
To achieve spectral alignment, we design the prompt as a graph $\mathcal{G}_{p}=(\mathcal{V}_{p},\mathbf{P},\mathcal{E}_{p})$. Here $\mathcal{V}_{p}$ is the prompt node set and $\mathbf{P}\in\mathbb{R}^{N_{p}\times d}$ are learnable node representations, where $N_{p}$ is the number of prompt nodes and $d$ is the feature dimension matching that of original graph nodes. Edges $\mathcal{E}_{p}$ are constructed based on the similarity between prompt nodes $e_{ij}=\mathbb{I}\big{(}{\sigma(\mathbf{p}_{i}\cdot\mathbf{p}_{j}^{T})>\tau_{inner}}\big{)}$, where $\mathbb{I}$ is the indicator function and $\tau_{inner}$ is the pre-defined threshold for inner edges within the prompt graph. $\mathbb{I}\big{(}{\sigma(\mathbf{p}_{i}\cdot\mathbf{p}_{j}^{T})>\tau_{inner}}\big{)}$ outputs 1 if $\sigma(\mathbf{p}_{i}\cdot\mathbf{p}_{j}^{T})$ is larger than $\tau_{inner}$; otherwise it outputs 0.
$\mathbf{p}_{i}^{\prime}=\frac{\mathbf{p}_{i}-\mu_{p}}{\sigma_{p}}\sigma_{o}+\mu_{o},$ (10)
where $\mu_{p},\sigma_{p}$, $\mu_{o},\sigma_{o}$ are the mean and standard deviation of the prompt and original graphs, respectively. We insert edges between nodes in $\mathcal{G}_{p}$ and $\mathcal{G}$ to attach the prompt graph into the whole original graph by $e_{ij}=\mathbb{I}\big{(}{\sigma(\mathbf{p}_{i}^{\prime}\cdot\mathbf{x}_{j}^{T})>\tau_{cross}}\big{)}$, where $\tau_{cross}$ denotes the pre-defined threshold for cross edges. Adjusting $\tau_{cross}$ controls node similarity, enabling selective introduction of diverse frequency components. The final prompted graph $\tilde{\mathcal{G}}=\psi(\mathcal{G},\mathcal{G}_{p})$, where $\psi$ denotes the insertion strategy. To establish the theoretical foundation of our spectral adaptation approach, Theorem 4.1 guarantees the universal adaptability of prompted graphs.
Theorem 4.1.
Given a pre-trained model $F$, along with two graphs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$. Assume their spectral distributions differ, i.e., $S_{high1}\neq S_{high2}$. Then, there exists a prompt graph $\mathcal{G}_{p}$ such that $F\big{(}\psi(\mathcal{G}_{1},\mathcal{G}_{p})\big{)}=F(\mathcal{G}_{2})$, where $\psi$ denotes the insertion strategy of $\mathcal{G}_{p}$.
Proof in Appendix A.3. Theorem 4.1 shows the prompted graph can theoretically equal graphs of any spectral distribution. Using Theorems 4.1 and 3.3, we get the following spectral alignment corollary:
@@ 4.3 Spectral-Aligned Prompt Tuning
The proof is in Appendix A.4. Corollary 4.2 shows using a proper prompt graph enhances performance by aligning the downstream graph’s spectral distribution with pre-trained spectral knowledge.
Due to pre-trained knowledge disparities among graph filters, we use a fine-grained spectral alignment method. For each graph filter $g_{k,C-k}$, we use a dedicated prompt graph $\mathcal{G}_{p}^{k}=(\mathcal{V}_{p}^{k},\mathbf{P}^{k},\mathcal{E}_{p}^{k})$, resulting $\tilde{\mathcal{G}}^{k}=\psi(\mathcal{G},\mathcal{G}_{p}^{k})$. Then the representations are generated and integrated as in the pre-trained stage:
$\mathbf{\tilde{Z}}^{k}=MLP\Big{(}g_{k,C-k}(\tilde{\mathbf{L}}^{k})\mathbf{\tilde{X}}^{k})\Big{)},\quad\mathbf{\tilde{z}}_{i}=\sum_{k=0}^{C}\frac{\exp(\mathbf{w}^{k})}{\sum_{j=0}^{C}\exp(\mathbf{w}^{j})}\odot\mathbf{\tilde{z}}_{i}^{k},$ (11)
where $\tilde{\mathbf{L}}^{k}$ and $\mathbf{\tilde{x}}_{i}^{k}$ are the Laplacian matrix and node features in the prompted graph $\tilde{\mathcal{G}}^{k}$. Here both MLP and $\mathbf{w}^{k}$ are the frozen pre-trained parameters. By aligning the downstream graph with the spectral characteristics of each filter, the model can fully utilize the pre-trained knowledge.
Model Optimization. The integrated representation $\tilde{\mathbf{z}}_{i}$ is fed into a one-layer MLP that acts as the classification task head. The model is optimized with cross-entropy loss:
$\mathbf{l}_{i}=Softmax\big{(}MLP(\tilde{\mathbf{z}}_{i})\big{)},$ (12)
$\mathcal{L}_{down}(\mathbf{\omega};\mathbf{\theta}^{*})=-\frac{1}{N}\sum_{i=1}^{N}\log(\mathbf{l}_{i})\mathbf{y}_{i}^{T},$ (13)
where $\mathbf{y}_{i}$ denotes the one-hot label vector. The frozen pre-trained parameter $\mathbf{\theta}^{*}$ includes the MLP layer in each graph filter and integration weight vector $\mathbf{w}^{k}$, the optimized parameter $\mathbf{\omega}$ includes the parameters of prompt nodes $\mathbf{P}^{k}$ and the MLP layer in the task head.
We discuss the parameter and time complexity of our model in Appendix C.3. Compared with a $(C+1)$-layer GCN, where $C+1$ is the number of graph filters in our backbone, we adopt a lightweight tuning and only introduce very limited additional parameters.

### v3

@@ 3.2.1 Spectral Distribution Diversity
3.2.2 Spectral Specificity and Generalization Limitation
Having established the spectral distribution diversity, we now investigate how this diversity impacts GNN generalization. We aim to prove the Spectral Specificity principle: a graph filter performs optimally only when its spectral response matches the input graph’s spectrum.
We employ the Spectral Regression Loss (SRL) Lei et al. (2022) to evaluate the filter $g$. Minimizing SRL serves as a proxy for maximizing downstream performance.
@@ 3.2.2 Spectral Specificity and Generalization Limitation
Given a pre-trained graph filter $g$ and an input graph $\mathcal{G}$ with spectral signal energy $\mathbf{\hat{x}}=(\hat{x}_{1},\ldots,\hat{x}_{N})^{T}$, to minimize the upper bound of the SRL, the filter response $g(\lambda_{i})$ must be positively correlated with the signal energy $\hat{x}_{i}^{2}$. Specifically, the filter should assign larger amplification $g(\lambda_{i})$ to frequencies where the signal energy $\hat{x}_{i}^{2}$ is larger.
The detail of SRL and proof are provided in Appendix A.2. Theorem 3.3 confirms that optimal performance requires alignment between pre-trained spectral filters and the downstream graph’s intrinsic spectrum. We verify the claim by testing spectral graph filters (i.e., low/mid/high-pass) on CSBM datasets from Figure 1(a). Results (details in Appendix B.5) show distinct graph filters excel on different graph types, aligning with our theorem. This imposes two fundamental requirements.
First, the pre-trained spectral knowledge must be abundant. Since downstream spectral distributions vary widely, a single-filter backbone suffers from irreversible information loss when signals fall outside its fixed passband. Thus, a hybrid spectral backbone is essential to ensure an abundant knowledge basis.
Second, and more crucially, this abundant knowledge must be fully utilized. Since each pre-trained filter captures a distinct frequency band, a specific downstream graph typically aligns with only a limited subset of filters, leaving the rich knowledge encoded in the remaining non-aligned filters underutilized. This dilemma motivates our method, which employs prompt tuning to actively transform the input spectrum, ensuring the abundant frozen spectral knowledge is fully utilized.
@@ 4 Methodology
Based on the theoretical analysis in Section 3.2, we propose HS-GPPT. The framework consists of two stages: (1) Hybrid Spectral Pre-training, where we construct a backbone (Section 4.1) to capture comprehensive spectral knowledge (Section 4.2); and (2) Spectral-Aligned Prompt Tuning, where we align downstream graphs with the pre-trained backbone to enhance the knowledge transfer (Section 4.3). The overall framework is in Figure 2, Algorithm 1, and Algorithm 2 in the appendix.
@@ 4.1 Hybrid Spectral Filter Backbone
To generalize across graphs with varying spectral distributions, we require a backbone capable of capturing information from a wide range of frequencies. We adopt the Beta Wavelet GNN (BWGNN) Tang et al. (2022) as our backbone implementation due to its flexible band-pass properties, where each focuses on different frequency components. The graph filters are formulated as:
$\small g_{k,r}(\mathbf{L})=\frac{{\frac{\mathbf{L}}{2}}^{k}(\mathbf{I}-\frac{\mathbf{L}}{2})^{r}}{2B(k+1,r+1)},$ (4)
where $B(k+1,r+1)=\frac{k!r!}{(k+r+1)!}$ is the Beta constant. With different constant values of $k$ and $r$, $g_{k,r}$ possesses different spectral characteristics. By setting $k+r=C$ as a constant, we obtain a group of filters $g_{\beta}=(g_{0,C},g_{1,C-1},\cdots,g_{C,0})$. Here, $g_{k,C-k}$ acts as a specific spectral filter. More details and visualization of BWGNN are in Appendix C.1 and Figure 6. We also evaluate the performance of other hybrid GNN backbones in Section 5.4. With these filters, we extract spectral representations from multiple filters:
$\mathbf{Z}^{k}=\text{MLP}_{k}\Big(g_{k,C-k}(\mathbf{L})\mathbf{X}\Big),\quad k\in\{0,\dots,C\}.$ (5)
@@ 4.2 Hybrid Spectral Pre-training
The hybrid backbone generates diverse spectral representations $\mathbf{Z}^{k}$. To initialize these filters and learn their effective integration, we employ self-supervised contrastive learning. Building on previous work Chen et al. (2024a), which uses local-global contrastive learning Velickovic et al. (2019) to align low- and high-frequency views, we extend this to multiple spectral views. We perform contrastive learning between node embeddings in different views (as local patches) and integrated graph embeddings (as global summaries), as shown in the upper part of Figure 2.
@@
To derive the global summary, we first integrate node representations from different spectral views. Since the spectral distribution diversity across feature dimensions, we employ an element-wise weighted sum to fuse the views adaptively:
$\small\mathbf{z}_{i}=\sum_{k=0}^{C}\frac{\exp(\mathbf{w}^{k})}{\sum_{j=0}^{C}\exp(\mathbf{w}^{j})}\odot\mathbf{z}_{i}^{k},$ (6)
@@
where $\odot$ denotes the element-wise product and $\mathbf{w}^{k}$ is the learnable integration weight vector. We then apply mean pooling over all nodes to obtain the global graph summary $\mathbf{z_{g}}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{z}_{i}$.
@@
For the contrastive objective, we generate negative samples by corrupting the graph structure (i.e., shuffling nodes while preserving the adjacency matrix), yielding negative embeddings $\mathbf{z}_{i}^{k-}$. We use a discriminator $\mathcal{D}$ to measure the agreement between node and graph embeddings $\mathcal{D}(\mathbf{z}_{i}^{k},\mathbf{z}_{g})=\sigma(\mathbf{z}_{i}^{k}\mathbf{W}\mathbf{z}_{g}^{T})\in[0,1]$, where $\mathbf{W}\in\mathbb{R}^{d\times d}$ denotes the weight matrix and $\sigma$ denotes the sigmoid activation function. The multi-view local-global mutual information maximization objective is formulated as:
$\small\mathcal{L}_{pre}(\mathbf{\theta})=-\frac{1}{CN}\sum_{k=0}^{C}\sum_{i=1}^{N}\Big(\log\mathcal{D}(\mathbf{z}_{i}^{k},\mathbf{z}_{g})+\log(1-\mathcal{D}(\mathbf{z}_{i}^{k-},\mathbf{z}_{g}))\Big).$ (7)
The learnable parameters $\mathbf{\theta}$ include MLP parameters in each graph filter, integration weight vectors $\mathbf{w}^{k}$ for k=(0,1,,C)k=(0,1,...,C), and the weight matrix of discriminator $\mathcal{D}$. Through this process, the model captures comprehensive spectral knowledge and learns to integrate it effectively. After pre-training, these parameters are fixed to serve as a stable spectral basis for the subsequent prompt tuning.
4.3 Spectral-Aligned Prompt Tuning
During the pre-training stage, the backbone learns a diverse set of graph filters and their integration schemes, learning rich spectral knowledge across different frequency bands. However, a specific downstream graph often exhibits a specific spectral bandwidth. According to the Spectral Specificity principle (Theorem 3.3), this mismatch leads to knowledge underutilization: only the filters matching the graph’s dominant frequency are fully utilized, while other valuable filters remain underutilized.
To address this, we propose Spectral-Aligned Prompt Tuning. Instead of modifying the frozen model, we optimize the input graph structure to align with each pre-trained filter individually. By transforming the graph’s spectral distribution to match the spectrum of every filter, we ensure that all pre-trained spectral knowledge is effectively activated, maximizing knowledge transfer (see Figure 2, lower part).
Prompt Construction. We design the prompt as a graph $\mathcal{G}_{p}=(\mathcal{V}_{p},\mathbf{P},\mathcal{E}_{p})$. Here $\mathcal{V}_{p}$ is the prompt node set and $\mathbf{P}\in\mathbb{R}^{N_{p}\times d}$ are learnable node representations, where $N_{p}$ is the number of prompt nodes and $d$ is the feature dimension matching that of original graph nodes. Edges $\mathcal{E}_{p}$ are constructed based on the similarity between prompt nodes $e_{ij}=\mathbb{I}\big({\sigma(\mathbf{p}_{i}\cdot\mathbf{p}_{j}^{T})>\tau_{inner}}\big)$, where $\mathbb{I}$ is the indicator function and $\tau_{inner}$ is the pre-defined threshold for inner edges within the prompt graph. $\mathbb{I}\big({\sigma(\mathbf{p}_{i}\cdot\mathbf{p}_{j}^{T})>\tau_{inner}}\big)$ outputs 1 if $\sigma(\mathbf{p}_{i}\cdot\mathbf{p}_{j}^{T})$ is larger than $\tau_{inner}$; otherwise it outputs 0.
$\small\mathbf{p}_{i}^{\prime}=\frac{\mathbf{p}_{i}-\mu_{p}}{\sigma_{p}}\sigma_{o}+\mu_{o},$ (8)
where $\mu_{p},\sigma_{p}$, $\mu_{o},\sigma_{o}$ are the mean and standard deviation of the prompt and original graphs, respectively. We insert edges between nodes in $\mathcal{G}_{p}$ and $\mathcal{G}$ to attach the prompt graph into the whole original graph by $e_{ij}=\mathbb{I}\big({\sigma(\mathbf{p}_{i}^{\prime}\cdot\mathbf{x}_{j}^{T})>\tau_{cross}}\big)$, where $\tau_{cross}$ denotes the pre-defined threshold for cross edges. Adjusting $\tau_{cross}$ controls node similarity, enabling selective introduction of diverse frequency components. The final prompted graph $\tilde{\mathcal{G}}=\psi(\mathcal{G},\mathcal{G}_{p})$, where $\psi$ denotes the insertion strategy.
@@ 4.3 Spectral-Aligned Prompt Tuning
Table 1: Performance comparison on 5-shot node classification under the transductive setting. Bold signifies the best result across all methods, while underline highlights the best baseline result. Additional results (e.g., standard deviations, baselines with other backbones) are in Appendix E.3 and E.4.
@@
Table 2: Performance comparison on 5-Shot node classification under the inductive setting. Here ‘Source’ denotes the pre-training datasets and ‘Target’ denotes the downstream dataset. More results are in Appendix E.3 and E.4.
Theoretical Guarantee. To theoretically justify our approach, we provide Theorem 4.1, which guarantees the universal adaptability of the prompted graph.
Theorem 4.1 (Spectral Adaptability).
Given a pre-trained model $F$, along with two graphs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, assume their spectral distributions differ, i.e., $S_{high1}\neq S_{high2}$. Then, there exists a prompt graph $\mathcal{G}_{p}$ such that $F\big(\psi(\mathcal{G}_{1},\mathcal{G}_{p})\big)=F(\mathcal{G}_{2})$, where $\psi$ denotes the insertion strategy.
Proof in Appendix A.3. Theorem 4.1 indicates that the prompted graph can theoretically approximate graphs of any spectral distribution. Combining this capability with the specificity requirement (Theorem 3.3), we derive the following alignment guarantee:
@@
The proof is in Appendix A.4. Corollary 4.2 confirms that using a proper prompt enhances performance by aligning the downstream graph’s spectrum with pre-trained knowledge.
Filter-Specific Alignment Strategy. To fully utilize the backbone, we cannot rely on a single global prompt, as it might fail to align with the disjoint passbands of the $C+1$ filters simultaneously. Therefore, we employ a fine-grained filter-specific alignment strategy. For each graph filter $g_{k,C-k}$ in the backbone, we assign a dedicated prompt graph $\mathcal{G}_{p}^{k}=(\mathcal{V}_{p}^{k},\mathbf{P}^{k},\mathcal{E}_{p}^{k})$. This generates independent prompted graphs $\tilde{\mathcal{G}}^{k}=\psi(\mathcal{G},\mathcal{G}_{p}^{k})$, each optimized to activate a specific spectral view. The representations are then computed and integrated:
$\small\mathbf{\tilde{Z}}^{k}=\text{MLP}_{k}\Big(g_{k,C-k}(\tilde{\mathbf{L}}^{k})\mathbf{\tilde{X}}^{k}\Big),\quad\mathbf{\tilde{z}}_{i}=\sum_{k=0}^{C}\frac{\exp(\mathbf{w}^{k})}{\sum_{j=0}^{C}\exp(\mathbf{w}^{j})}\odot\mathbf{\tilde{z}}_{i}^{k},$ (9)
where $\tilde{\mathbf{L}}^{k}$ and $\mathbf{\tilde{x}}_{i}^{k}$ are the Laplacian matrix and node features in the prompted graph $\tilde{\mathcal{G}}^{k}$. Here both MLP and $\mathbf{w}^{k}$ are the frozen pre-trained parameters. By aligning the downstream graph with the specific spectral characteristics of each filter, the model ensures that all pre-trained knowledge is fully used.
Model Optimization. The integrated representation $\tilde{\mathbf{z}}_{i}$ is fed into a lightweight task head (one-layer MLP) for classification. The model is optimized via the cross-entropy loss:
$\small\mathbf{l}_{i}=\text{Softmax}\big(\text{MLP}_{head}(\tilde{\mathbf{z}}_{i})\big),$ (10)
$\mathcal{L}_{down}(\omega;\theta^{*})=-\frac{1}{N}\sum_{i=1}^{N}\log(\mathbf{l}_{i})\mathbf{y}_{i}^{T},$ (11)
where $\mathbf{y}_{i}$ is the one-hot label vector. Note that $\theta^{*}$ represents the frozen pre-trained parameters (including filter MLPs and integration weights $\mathbf{w}^{k}$), while $\omega$ denotes the learnable parameters (prompt features $\mathbf{P}^{k}$ and task head).
We discuss the parameter and time complexity in Appendix C.3. Compared with a $(C+1)$-layer GCN, where $C+1$ is the number of graph filters in our backbone, our method introduces minimal additional parameters, ensuring efficiency.